OpenStudy (shayaan_mustafa):

Show that \[\lim_{x \rightarrow 0^{+}}(1/x - 1/x ^{2}) = -\infty\]

OpenStudy (angela210793):

An idea... I think u get infinit-infinit= undefined form so u must use L'Hopital's rule....

OpenStudy (shayaan_mustafa):

yes inf-inf=inf. but how -inf?

OpenStudy (angela210793):

no inf-inf is undefined

OpenStudy (shayaan_mustafa):

I know. madam but in question it must be -inf.

OpenStudy (shayaan_mustafa):

hey. anyone help me with this?

OpenStudy (anonymous):

subtract

OpenStudy (shayaan_mustafa):

@satellite73 help me please

OpenStudy (shayaan_mustafa):

yeah. but how to get -inf?

OpenStudy (anonymous):

\[\frac{1}{x}-\frac{1}{x^2}=\frac{x-1}{x^2}\] now take the limit

OpenStudy (shayaan_mustafa):

it will go to infinity.

OpenStudy (anonymous):

x is going to zero, so denominator is going to zero and numerator is going to -1 denominator is always positive so you go to minus infinity

OpenStudy (shayaan_mustafa):

but one thing. if I take lim at this step i.e. 1/x - 1/x^2 Then?

OpenStudy (anonymous):

you cannot take the limit in that step by taking the limits separately you have to subtract first if you try to take the limit one piece at the time you get \(\infty-\infty\) which is undetermined

OpenStudy (shayaan_mustafa):

oh .. that's fine thanx buddy.

OpenStudy (shayaan_mustafa):

but according to howard anton calculus we can take limit separately Theorem#2.2.2 6th edition a new horizon.

OpenStudy (shayaan_mustafa):

@satellite73